Cost/King-size Set

Raw Cotton

 

$28.00

Spinning/Weaving/Dyeing

 

$12.00

Cut/Sew/Finishing

 

$10.00

Material Transportation

 

$ 3.00

Factory Fee

 

$16.00

Inspection and Import Fees

 

$14.00

Ocean Freight/Insurance

 

$ 15.00

Warehousing

 

$ 8.00

Packaging

 

$15.00

Promotion

 

$30.00

Customer Shipping

 

$15.00

Given the cost per king-size sheet set above, and assuming the manufacturer has total fixed costs of $600,000 and estimates first year sales will be 50,000 sets

Determine the price to consumers if the company desires a 50 percent mark-upon cost.

 

1
Expert's answer
2021-06-22T04:05:01-0400

The variable cost/king-size set is:

28+12+10+3+16+14+15+8+15+30+15 = $166

If total fixed costs = $600,000 and estimated sales = 50,000 sets

The Fixed cost/king-size set is: 600,000/50,000 = $12

Cost price/unit = variable cost per unit + fixed cost per unit

Cost price per king-size set is therefore : 166 + 12 = $178


Markup percentage = [(selling price/unit - cost price/unit)/cost price per unit] * 100

Therefore: 50 = [(selling price/unit - 178)/178] * 100

50 * 178/100 = Selling price/unit - 178

selling price per unit = 89 + 178

selling price per unit = $267

The company should sell each king-size set at $267 in order to meet a 50% markup on cost.